# Artworks Reimagined: Exploring Human-AI Co-Creation through Body Prompting

### v1

@@ 3.3. Study Design and Procedure
Architecture.
The study was conducted in accordance with the policies of the University of Jyväskylä’s Institutional Review Board (IRB) and the Finnish National Board on Research Integrity (TENK). Children under the age of 18 were allowed to use the installation under the supervision of adults, but were excluded from the interviews. In this article, we only focus on the data collected from 79 adults who agreed to an interview.
3.4. Data Collection and Analysis
We developed an interview guide consisting of 12 questions. Besides basic demographics, the research assistants asked participants about their motivation for selecting the artwork, the body prompt, and the choice of private or public posing area. The interviewers also inquired how participants experienced body prompting and what they wanted to express with their body prompt. The structured questions were followed by open-ended follow-up questions (“can you describe why?”). A Likert-scale question captured how enjoyable the experience was for the participant (from 1 – Not At All Pleasant to 5 – Extremely Pleasant). The interview also included the ten-item questionnaire on key personality traits (commonly referred to as “Big Five”) (Digman, 1990).
The generated image was revealed on the screen at the end of the interview and participants were asked whether the image expressed their original intent, thus giving insights into co-creative aspects of body prompting, and any other thoughts about their body prompt and the generated image.
The interview data was transcribed using a text-to-speech service, translated from Finnish into English using Google Translate, and manually reviewed and corrected by three research assistants who listened to the recordings and manually corrected translation errors. The qualitative data was analyzed following a grounded theory approach (Glaser and Strauss, 1967). Three authors (one Postdoc in Cognitive Science, one PostDoc in Computer Science, and one PhD student with professional experience in user experience design) discussed a preliminary coding scheme for each question as a starting point for inductive coding. The three authors then individually coded the responses of a sample of 30 participants. In a discussion following this individual coding session, the authors noticed that coding the data was straightforward in all cases (McDonald et al., 2019). That means, in all but a very few cases, it was easy to identify themes in the data, and in two other questionnaire items, the responses were too diverse to identify common themes. Therefore, it was decided that one author would inductively code the interview data.
The generated images and the body prompts (i.e., poses) were rated by the three authors along three criteria: a) dynamism of the source artwork (low/medium/high), b) dynamism of the participant’s pose (low/medium/high), and c) a comparison of the source artwork and generated image in terms of whether a change in narrative took place. For a) and b), the authors collaboratively agreed on a codebook, as listed in Table 2 in Appendix A. With the metric of ‘dynamism’, we intend to capture the expressiveness of the artwork and body prompt. A static pose was rated to be of low expressiveness, whereas a highly dynamic pose was rated as being very expressive. For c), the authors agreed to code a change in narrative if a subject engaged in a different activity or had a different expression (e.g., static versus humorous), or if the model hallucinated artifacts that caused or strongly contributed to a change in narrative. After a first rating round, the ratings were discussed and each author revised their ratings for internal consistency.
Employing Fleiss’ Kappa for inter-rater reliability assessment, our study found high agreement levels among the three raters across all three variables: source artwork dynamism ($\kappa=0.96$), pose dynamism ($\kappa=0.86$), and narrative change ($\kappa=0.88$). These results underscore the raters’ cohesive interpretations and a consistent and robust evaluation process for the studied variables. After calculating inter-rater agreement, a majority vote between the three authors was used to produce the final set of ratings.
In addition, one author analyzed and coded the difference between the source artwork and the body prompt, characterizing the latter as either being an imitation (re-creation) of the source artwork or an attempt to ‘reimagine’ (reinterpret) the artwork with a body prompt that significantly differs from the source artwork. The latter equates to wanting to change the narrative between source artwork and generated artwork via the body prompt. It was further coded whether this narrative change in the resulting image originated from the generative model (e.g., by hallucinating artifacts or facial expressions) or the user (via the body prompt), with options ‘model’, ‘user’, ‘both’, and ‘n/a’ (meaning no major change in narrative took place).
3.5. Technical Implementation
We developed a software architecture based on T2I-Adapter (Mou et al., 2023) with four different inputs (depicted in Figure 3):
  1. (1)
    Pose: The pose was detected from the body prompt photo and used to guide the image generation with openpose-based ControlNet (Zhang et al., 2023; Cao et al., 2021).
  2. (2)
    Artwork: Style information was extracted from the selected source artwork and used as an input for the T2I-Adapter.
  3. (3)
    Prompt: A textual prompt was generated from the selected source artwork using CLIP-Interrogator (version 0.6.0). [^4] https://github.com/pharmapsychotic/clip-interrogator
  4. (4)
    Negative prompt: A negative textual prompt was manually designed, based on best practices in text-to-image generation (Oppenlaender, 2023). This negative prompt was used for all image generations to improve the quality of outputs and prevent unsafe generations (Qu et al., 2023).
Images were generated with Stable Diffusion v1.5 (Rombach et al., 2022) with 50 steps and cfg 8.0. The generated images were upscaled twice with Real-ESRGAN for viewing on the large public display. GFPGAN was used to enhance faces. The image generation architecture was packaged as a cog container and uploaded to Replicate.com where it is publicly available for inference via an API. [^5] https://replicate.com/joetm/camerabooth-openpose-style
Two React-based web applications were developed, one for body prompting and one for viewing the results. Both applications were hosted in AWS S3 buckets, using an ssl-encrypted AWS Cloudfront to provide a secure connection.
The first application guided users through the process with a linear application flow (c.f. Figure 1). The application captured the photo from the webcam MediaStream using the MediaDevices API once the timer hit zero. An AWS Gateway was set up to proxy requests from the web application to Replicate’s API. Results from the API were received with a webhook (AWS Lambda function) and stored in a JSON file hosted in an S3 bucket. The application was automatically reset after 60 seconds once the final application screen was reached. The main purpose of this timed reset was to reduce queuing for participants at the public viewing area — the bottleneck of our study design — and to give more time to the subsequent interview. The source artworks were shuffled after each image generation.
The second application enabled the research assistants to scroll through the generated images with keyboard arrow keys. The application used long-polling to fetch the list of generated images. A red square on the screen indicated the number of newly generated images available for viewing. The participants pose was shown in the corner of the screen. Generated images were viewed either on a public display or in private on two laptops with privacy screens, depending on which posing area was chosen by the participant.
[…]
Example of body prompts and resulting generated images.
4. Results
This section first provides information on the participants and then proceeds to answer our research questions.
4.1. Participants
Interviews were conducted with 79 participants after they used the installation. Participants were on average 37 years old ($Median=35$, $SD=13.5$, $Max=68$) and included 46 women, 32 men, and one non-binary participant.
Due to the study being conducted in the field without control over the number of participants per condition (i.e., group versus solo participation, body prompting strategy, artwork choice, and public versus private posing area), we focus on describing the images in Section 4.2 and qualitatively analyzing the interview data in sections 4.34.6. For a rich description, we contextualize participant quotes with the codes listed in Table 1.
Table 1. Contextual information on participants.
@@ 4.1. Participants
For participating groups, interviews were held with one group member, but other group members were allowed to comment. For simplicity, we refer to both groups and single participants as ‘participant’ in this work. In the following section, we describe the images generated at the event.
4.2. RQ1: Generated Images
In total, 172 images were created during the six-hour event (112 images in public and 60 images in the private booth). Due to the diversity of source artworks, the body prompts are difficult to compare. In this section, we provide a brief descriptive overview of the 172 generated images and body prompts. This section mainly reports on general trends, while specific outstanding instances are highlighted in Section 4.6.1. A gallery of all generated images is available at https://artworksreimagined.com.
@@ 4.2. RQ1: Generated Images
The body prompts had varying degrees of expressiveness. Some participants simply stood motionless in front of the camera in a neutral pose with arms dangling beside their hips. Other participants were more dynamic and expressive, raising limbs into the air. A few participants went to great lengths to produce their body prompts (as depicted in Figure 5, bottom center). The generative model’s output would reflect the same number of posing subjects. Overall, there were no noticeable differences between the body prompts performed in public and private, with a few exception (described in Section 4.6.1).
We classified whether a change in narrative had taken place between source artwork and generated image. In almost half of the generated images (46.8%) a change originated from the participant’s body prompt, which means the pose introduced significant modifications in the generated image. In another 29.1% of the generated images, no change in narrative took place and the new generated image appeared to be a reproduction of the source artwork’s original idea, with some minor changes (e.g., number of subjects).
About a quarter of the participants (24.1%) encountered surprising generative artifacts in the generated image. Co-creation with AI can involve surprising hallucinations by the generative model. With hallucinations, we specifically refer to artifacts introduced by the generative model that cannot be found in the source artwork or in the participant’s body prompt, often resulting in a change in narrative between the source artwork and the generated image. In this study, hallucinatory artifacts included door frames, wooden sticks, fields with houses in the distances, animals, and more. However, these generated elements in the image did not necessarily deviate from the narrative portrayed in the source artwork.
About 8% of the participants encountered surprising artifacts originating from the generative model that clearly changed the narrative. Another 16.2% of participants encountered images where the narrative change originated both from the model and the participant. These were often images where the participant’s pose deviated from the source artwork, with the AI adding surprising elements, such as additional characters or changes to the subjects. In other instances, the model added minor hallucinatory elements that still resulted in a change of narrative. For instance, in the generated image in Figure 5, bottom right, the model appears to have added COVID face masks which are not present in the source artwork. In another funny example, the generative model seemingly changed the gender of a subject from a blonde woman to bald man (see Figure 5, middle row, center).
@@
Figure 5. Examples of a change in narrative originating from the user (top row), the generative model (middle row), and both (bottom row), with the respective source artwork and body prompt. Hallucinations from the model included, for instance, additional characters in the generated images (middle row, right, and bottom row, left), but also more subtle changes, such as a change in gender of a subject (middle row, center), face masks (bottom row, right), or a man walking an apple on a leash (middle row, left).
@@
4.3. RQ2: User Experience
4.3.1. Experiencing body prompting
Participants generally found the experience of body prompting highly pleasant ($Mean=4.58$ on a Likert-scale from 1 to 5, $SD=0.62$). Creating images via body prompting was deemed to be overall a great experience ($Mean=4.58$, $SD=0.65$). When asked to describe why, participants provided different explanations. The most common reasons were that participants found it fun, interesting, and easy. The ease of the interaction via body prompting was commented on often. The in-app instructions were easy to understand and follow, providing a “fluent” (A38, D70) and “effortless” (C43) user experience.
@@ 4.3.1. Experiencing body prompting
Body prompting was perceived as fun and lighthearted, with “not too much pressure participating” (A35). Many participants called it a “fun experiment” (e.g., C28, G79) and some participants appreciated that “you get to be creative and go crazy” (C28). Many participants wanted to try the installation a second time and experiment with different body prompts, such as F62 who wanted to “try a completely different pose than in the artwork”. Speaking about Hokusai’s “The Great Wave off Kanagawa”, H30 commented they “wanted to see if AI transforms us into humans or will it convert us to waves as there were no humans in the original artwork”. As depicted in Figure 5 (top right), the AI indeed added two humans in this case, although in other cases, it did not.
A small minority of participants voiced some discomfort and concerns about body prompting. Most of these comments, however, related to the study design and the technical limitations of the installation. For instance, three participants commented on the wait time which was a fixed part of the study design. Some participants commented on the countdown timer. On the one hand, F57 thought “it was quite a long time to stand and look at your own picture”. On the other hand, some participants wanted to have more time to decide on their body prompt. G66, for instance, thought the picture was taken too quickly and “there was no time to think”. Addressing another technical limitation, A24 thought it was a fun experience, but voiced disappointment about the facial expression not being captured by the body prompt. While privacy-by-design was explicitly appreciated by two participants (G61 and B65), a few other participants still voiced concerns about being photographed. This included comments from E18 and G78 who categorically do not like being photographed. G78 mentioned the picture-taking being “a little bit distressing”. A2 thought it was daring to perform in public, and C43 voiced concerns about being watched: “When I started to do my pose I started to think is there someone watching. Other than that it was fun to see what is coming up”.
We could observe that body prompting by a group of people generally tended to be more entertaining, both for the participants performing the body prompting and the audience. Groups of participants developed their own dynamics leading to playful interactions with the image generation system. Selecting the right source artwork often involved short discussions in which participants agreed on their body prompt. These group negotiations could, obviously, not be observed in single participants. However, some participants instead asked the audience on what body prompt to perform, turning the audience from passive spectators to active participants.
@@
4.3.2. Experiencing AI co-creation via body prompting
@@
In this section, we describe how participants experienced the co-creative embodied interaction with the generative AI. Note that this experience is, in our study, influenced by the pre-selected source artwork and not every combination of source artwork and body prompt worked equally well. The participants’ experiences are, therefore, diverse and difficult to summarize. Experiences with AI co-creation in our study can broadly be divided into whether the generated image met (or exceeded) the participants’ expectations or not. Because the latter category is more insightful due to friction in the experience, we place emphasis on this latter category.
@@
Among the participants who stated their expectations were met ($n=25$), participants commented the generated image was as they had imagined. C3, for instance, said the generated imagine “expresses [the body prompt] pretty well”. Some participants found the image better than expected, often commenting on the funniness of the image. In this participant group, the generated image matched the dynamism of the body prompt. Participants in this group could recognize themselves and described the experience as interesting and fun.
@@
For the participants who encountered some friction ($n=54$), the installation was also fun and interesting and overall a very good experience, with some limitations. Participants often commented on the installation’s technical limitations, including its inability to reproduce facial expressions and exact hand gestures. This left room for the generative model to add its own interpretation. A23 said that “the first thing that catches [her] eye is those hands, they somehow look a bit unnatural”. In other cases, issues with distorted body parts were even more pronounced, as A35 remarked, “the fingers look like spaghetti” and B68 exclaimed “look at your head, it’s like an alien!”. A24 also encountered issues with the facial expressions, commenting that “the face is completely distorted”. Some generated images had an uncanniness and strangeness that is often associated with AI-generated visuals. For some participants, the weird generative artifacts made the image “pleasant to look at but hard to relate” (A16). On the other hand, the uncanniness was positive for some participants, such as A24, who, despite the distorted face, thought “the face is amazing”, “better than [we] thought”, and “a positive surprise”. Many participants had an expectation that the system would produce their likeness. C1, for instance, commented “I expected that the face would resemble my own face”. This was even more pronounced in children, as discussed in Section 4.6.2.
@@
For a few participants, the uncanny artifacts shifted the mood of the image, sometimes making it more gloomy, sinister, or eerie than anticipated. For instance, after seeing the unnatural hands, A35 said “there is something disturbing” and A33 commented “it’s scarier than we thought”. A17 thought there was “a more serious atmosphere” in the generated image. Some participants said this mood change was unexpected, with a few participants commenting that they had wanted the image to be “more cheerful” (F62). B67 commented, after finding the outcome “sad and gloomy”, that the generated image “does not reflect what we meant”. A13 also commented that “it was supposed to be funny, and this is more grim”. On the other hand, one participant (A5) thought the generated image had a nice “Christmas feeling” that reminded him of a “Christmas carol”.
@@
Participants generally found it difficult to understand why the generative model did what it did (even though the body prompt pose was presented to them). One participant tried to explain the mismatch between their likeness and the generated subject as being a mixture of characteristics of the source artwork and body prompt: “it has probably taken characteristics of both the original and me, so now it’s fun to compare the original to this” (F74). Most participants could not articulate reasons for deviations in the generated image from their body prompt or the source artwork, and very few tried to verbalize such explanations. This is partly due to many participants approaching the installation in a casual interaction without strong expectations. Participants also had difficulty imagining what the generated image would look like. E12 commented on the strong differential between expectation and outcome. According to her, “it has very hard to imagine how the outcome would be. That creates very strong expectations”. Without any reference in mind, she commented that it “affected the way you see the picture” and she also mentioned that seeing the generated images from other people gave her a better reference point and an anchor for her own body-prompted image generation.
4.4. RQ3: Participant Behavior

### v2

@@ 5.3. Participants
For participating groups, interviews were held with one group member, but other group members were allowed to comment. For simplicity, we refer to both groups and single participants as ‘participant’ in this work.
5.4. Results
In this section, we describe the images generated at the event (Section 5.4.1) and how participants experience the installation and co-creation with the AI (Section 5.4.2). We further describe the key choices and behavior of participants and different body prompting strategies (Section 5.4.3). Last, we report on the differences between the public and private setting (Section 5.4.4) and miscellaneous observations (Section 5.4.5). This section mainly reports on general trends, while specific outstanding instances are highlighted in Section 5.4.5.
5.4.1. RQ1: Generated images
In total, 172 images were created during the six-hour event (112 images in public and 60 images in the private booth). Due to the diversity of source artworks, the body prompts are difficult to compare. In this section, we provide a brief descriptive overview of the 172 generated images and the body prompts. A gallery of all generated images is available at https://artworksreimagined.com.
[…]
Example of body prompts and resulting generated images.
@@ 5.4.1. RQ1: Generated images
The body prompts had varying degrees of expressiveness. Some participants simply stood motionless in front of the camera in a neutral pose with arms dangling beside their hips. Other participants were more dynamic and expressive, raising limbs into the air. A few participants went to great lengths to produce their body prompts (as depicted in Figure 5, bottom center). The system’s output accurately reproduced the number of posing subjects. Overall, there were no noticeable differences between the body prompts performed in public and private, with a few exception (described in Section 5.4.5).
We classified whether a change in narrative had taken place between source artwork and generated image. In almost half of the generated images (46.8%) a change originated from the participant’s body prompt, which means the pose introduced significant modifications to the source image, shifting the narrative. In another 29.1% of the generated images, no shift in narrative took place and the new generated image appeared to be a reproduction of the source artwork’s original idea, with only some minor changes.
About a quarter of the participants (24.1%) encountered surprising generative artifacts. Co-creation with AI can involve surprising hallucinations by the generative model. With hallucinations, we specifically refer to artifacts introduced by the generative model that cannot be found in the source artwork or in the participant’s body prompt, often resulting in a shift in narrative between the source artwork and the generated image. In our study, hallucinatory artifacts included door frames, wooden sticks, fields with houses in the distances, animals, and more. These artifacts appeared in the generated image without being specifically prompted or being part of the source artwork. However, these generated elements in the image did not necessarily deviate from the narrative portrayed in the source artwork. Instead, they sometimes complemented what was already in the source artwork, or they complemented the body prompt.
About 8% of the participants encountered surprising artifacts originating from the generative model that clearly shifted the narrative. Another 16.2% of participants encountered images where the shift in narrative originated both from the model and the participant. These were often images where the participant’s pose sbstantially deviated from the source artwork, with the AI adding surprising elements, such as additional characters or changes to the subjects. In other instances, the model added minor hallucinatory elements that still resulted in a change of narrative. For instance, in the generated image in Figure 5, bottom right, the model appears to have added COVID face masks to the subjects’ faces. These masks were not present in the source artwork. In another funny example, the generative model seemingly changed the gender of a subject from a blonde woman to bald man (see Figure 5, middle row, center).
@@
Figure 5. Examples of a change in narrative originating from the user (top row), the generative model (middle row), and both (bottom row), with the respective source artwork (top right) and body prompt (bottom right). Hallucinations from the model included, for instance, additional characters in the generated images (middle row, right, and bottom row, left), but also more subtle changes, such as a change in gender of a subject (middle row, center), face masks (bottom row, right), or a man walking an apple as if it were a dog on a leash (middle row, left).
@@
5.4.2. RQ2: User experience
This part of our investigation looks at how participants experienced body prompting and how they experienced co-creation with the generative AI.
Experiencing body prompting
Participants generally found the experience of body prompting highly pleasant ($Mean=4.58$ on a Likert-scale from 1 to 5, $SD=0.62$). Creating images via body prompting was deemed to be an overall great experience ($Mean=4.58$, $SD=0.65$). When asked to describe why, participants provided different explanations. The most common reasons were that participants found it fun, interesting, and easy. The ease of the interaction via body prompting was commented on often. The in-app instructions were easy to understand and follow, providing a “fluent” (A38, D70) and “effortless” (C43) user experience.
@@ 5.4.2. RQ2: User experience
Body prompting was perceived as fun and lighthearted, with “not too much pressure participating” (A35). Many participants called it a “fun experiment” (e.g., C28, G79) and some participants appreciated that “you get to be creative and go crazy” (C28). Many participants wanted to try the installation a second time and experiment with different body prompts, such as F62 who wanted to “try a completely different pose than in the artwork”. Speaking about Hokusai’s The Great Wave off Kanagawa, H30 commented they “wanted to see if AI transforms us into humans or will it convert us to waves as there were no humans in the original artwork”. As depicted in Figure 5 (top right), the AI indeed added two humans in this case, although in other cases, it did not.
A small minority of participants voiced some discomfort and concerns about body prompting. Most of these comments, however, related to the study design and the technical limitations of the installation. For instance, three participants commented on the wait time in between prompting and viewing the results which was a fixed part of our study design. Some participants commented on the countdown timer. On the one hand, F57 thought “it was quite a long time to stand and look at your own picture”. On the other hand, some participants wanted to have more time to decide on their body prompt. G66, for instance, thought the picture was taken too quickly and “there was no time to think”. Addressing another technical limitation, A24 thought it was a fun experience, but voiced disappointment about the facial expression not being captured by the body prompt. While our privacy-by-design approach was explicitly appreciated by two participants (G61 and B65), a few other participants still voiced concerns about being photographed. This included comments from E18 and G78 who categorically do not like being photographed. G78 mentioned the picture-taking being “a little bit distressing”. A2 thought it was daring to perform in public, and C43 voiced concerns about being watched: “When I started to do my pose I started to think is there someone watching. Other than that it was fun to see what is coming up”.
We could observe that body prompting by a group of people generally tended to be more entertaining, both for the participants performing the body prompting and the on-looking audience. Groups of participants developed their own dynamics leading to playful interactions with the image generation system. Selecting the right source artwork often involved short discussions in which participants agreed on their body prompt. These group negotiations could, obviously, not be observed in single participants. However, some single participants instead asked the audience on what body prompt to perform, turning the audience from passive spectators to active participants.
@@
Experiencing AI co-creation via body prompting
@@
In this section, we describe how participants experienced the co-creative embodied interaction with the generative AI. Note that this experience is, in our study, biased by the pre-selected source artwork. Not every combination of source artwork and body prompt worked equally well. The participants’ experiences are, therefore, diverse and difficult to summarize. Experiences with AI co-creation in our study can broadly be divided into whether the generated image met (or exceeded) the participants’ expectations or not. Because the latter category is more insightful due to friction in the experience, we place emphasis on the minority of participants in the latter category.
@@
Among the participants who stated their expectations were met ($n=25$), participants commented the generated image was as they had imagined. C3, for instance, said the generated imagine “expresses [the body prompt] pretty well”. Some participants found the image better than expected, often commenting on the funniness of the image. In this participant group, the generated image matched the dynamism of the body prompt. Participants in this group could recognize themselves in the generated image, and described the experience as interesting and fun.
@@
For the participants who encountered some friction ($n=54$), the installation was also fun and interesting and overall a very good experience, with some minor limitations. Participants often commented on the installation’s technical limitations, including its inability to reproduce facial expressions and exact hand gestures. This left room for the generative model to add its own interpretation. A23 said that “the first thing that catches [her] eye is those hands, they somehow look a bit unnatural”. In other cases, issues with distorted body parts were even more pronounced, as A35 remarked, “the fingers look like spaghetti” and B68 exclaimed “look at your head, it’s like an alien!”. A24 also encountered issues with the facial expressions, commenting that “the face is completely distorted”. Some generated images had an uncanniness and strangeness that is often associated with AI-generated images. For some participants, the weird generative artifacts made the image “pleasant to look at but hard to relate” (A16). On the other hand, the uncanniness was positive for some participants, such as A24, who, despite the distorted face, thought “the face is amazing”, “better than [we] thought”, and “a positive surprise”. Many participants had an expectation that the system would produce their likeness. C1, for instance, commented “I expected that the face would resemble my own face”. This was even more pronounced in children, as discussed in Section 5.4.5.
@@
For a few participants, the uncanny artifacts shifted the mood of the image, sometimes making it more gloomy, sinister, or eerie than anticipated. For instance, after seeing the unnatural hands, A35 said “there is something disturbing” and A33 commented “it’s scarier than we thought”. A17 thought there was “a more serious atmosphere” in the generated image. Some participants said this shift in mood was unexpected, with a few participants commenting that they had wanted the image to be “more cheerful” (F62). B67 commented, after finding the outcome “sad and gloomy”, that the generated image “does not reflect what we meant”. A13 also commented that “it was supposed to be funny, and this is more grim”. On the other hand, one participant (A5) thought the generated image had a nice “Christmas feeling” that reminded him of a “Christmas carol”.
@@
Participants generally found it difficult to understand why the generative model did what it did (even though the body prompt pose was presented to them). One participant tried to explain the mismatch between their likeness and the generated subject as being a mixture of characteristics of the source artwork and body prompt: “it has probably taken characteristics of both the original and me, so now it’s fun to compare the original to this” (F74). Most participants could not articulate reasons for deviations in the generated image from their body prompt or the source artwork, and very few tried to verbalize such explanations. This is partly due to many participants approaching the installation in a casual interaction without strong expectations. Participants also had difficulty imagining what the generated image would look like. E12 commented on the, in her case, strong difference between expectation and outcome. According to her, “it has very hard to imagine how the outcome would be. That creates very strong expectations”. Without any reference in mind, she commented that it “affected the way you see the picture” and she also mentioned that seeing the generated images from other people gave her a better reference point and an anchor for her own body-prompted image generation.
5.4.3. RQ3: Participant behavior